PROBING ROTATIONALLY EXCITED STATES OF CH_{5}^{+} WITH DIFFUSION MONTE CARLO

\underline{CHARLOTTE E. HINKLE}, ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Protonated methane has long proven to be a challenging system for both experimentalists and theoreticians. The essentially flat potential surface comprised of 120 equivalent minima, coupled with a very fluxional molecule, make CH_{5}^{+} a challenging system to study. Using Diffusion Monte Carlo we have had previous success studying vibrationally excited states of CH_{5}^{+}. Here we focus on modeling rotationally excited states using Diffusion Monte Carlo. Following our success with H_{3}O^{+} and D_{3}O^{+} we define our rotationally excited states by placing nodes at the zeros in the real rotational eigenstates of a symmetric top. We use this approach to analyze rotationally excited states of CH_{5}^{+} through use of Fixed Node Diffusion Monte Carlo. We use the results of these simulations to analyze the rotation/vibration mixing in rotationally excited states of CH_{5}^{+}.